OpenStudy (anonymous):

If a2 + b2 + c2 = 250 and ab + bc + ca = 3, then a + b + c is

OpenStudy (anonymous):

shayan use here the formula of (a+b+c)^2 this would help

OpenStudy (asnaseer):

follow suggestion by waheedahmed and answer should become abvious

OpenStudy (anonymous):

thanks a lot 16 ans

OpenStudy (anonymous):

or -16?

OpenStudy (anonymous):

Area of triangle, having sides as 8, 11 and the perimeter 32 is

OpenStudy (anonymous):

(a + b + c)^2= a^2 + b^2 + c^2 + 2(ab + bc + ca), since, a^2 + b^2 + c^2 = 250 and ab + bc + ca = 3, (a + b + c)^2 = 250 + 2(3) = 256 = 16^2 therefore: a + b + c = 16 ,.

OpenStudy (anonymous):

why not -16?

OpenStudy (anonymous):

Area of triangle, having sides as 8, 11 and the perimeter 32 is

OpenStudy (anonymous):

area is always taken positive shayanreloaded thats why...

OpenStudy (anonymous):

lol these are 2 dif qs

OpenStudy (anonymous):

i have given u the answer of above shayan

OpenStudy (anonymous):

not the second one

OpenStudy (anonymous):

for first -16 is possible right? and help wid 2nd

OpenStudy (anonymous):

you had asked why not -16 for that

OpenStudy (anonymous):

no area cant be -16 shayan

OpenStudy (anonymous):

this is fr 1st question which is not bout area

OpenStudy (anonymous):

ok my mistake both 16 and -16 are possible sorry about that shayan ...i hope u would understand

OpenStudy (anonymous):

yeah i do now 2nd ques lol

OpenStudy (anonymous):

do you know the formula for perimeter shayan the solution is in it

OpenStudy (anonymous):

hm add sides?

OpenStudy (anonymous):

3rd side is 13 i dont think i rem. herons formula

OpenStudy (anonymous):

after getting third side u will find s=a+b+c/2

OpenStudy (anonymous):

16 ok

OpenStudy (anonymous):

16 is the answer

OpenStudy (anonymous):

now use this A=squareroot(s(s-a)(s-b)(s-c)

OpenStudy (anonymous):

did u get the answer shayan

OpenStudy (anonymous):

43.8?